OpenStudy (anonymous):

volume of a rectangular prism is 6c^3+14c^2+4c, and the dimensions are monomial and binomial. What are the dimensions of the prism

OpenStudy (anonymous):

V=l*w*h you can re-write : 6c^3+14c^2+4c=c*(c+2)(6c+2) = V so, your dimensions are: c, (c+2) & (6c+2) Good?